# devoted to memory of Alexei Alexandrovich Starobinsky

### v12

@@ I Introduction
The $f({\cal R})$ theory demonstrates that the effective gravitational coupling $K$ (it is the inverse Newton constant, $K=1/16\pi G$) and the scalar curvature ${\cal R}$ are connected by equation $K=df/d{\cal R}$. This suggests that $K$ and ${\cal R}$ are the thermodynamically conjugate variables.Volovik2023a ; Volovik2023b This pair of the gravitational variables is similar to the pair of the electrodynamic variables, electric field ${\bf E}$ and electric induction ${\bf D}$, which participate in the thermodynamics of dielectrics. Using the local thermodynamics with $T=H/\pi$, we obtained the general result for the total entropy of the Hubble volume, $S_{\rm bulk}=sV_{H}=4\pi KA=S_{\rm hor}$, where $K=df/d{\cal R}$ is the effective gravitational coupling.
@@
In Sec. II we consider the elements of the local thermodynamics of the de Sitter state. It is based on the observation that the de Sitter vacuum serves as the thermal bath for the external matter – the matter immersed into the de Sitter environment. This is different from the traditional consideration of the vacua of the quantum fields in the de Sitter spacetime, which uses the Euclidean action method. Example of the influence of the de Sitter vacuum to the external matter is provided by an atom in the de Sitter environment (Sec. II.1). As distinct from the atom in the flat space, the atom in the de Sitter vacuum has a certain probability of ionization. The rate of ionization is similar to the rate of ionization in the presence of the thermal bath with temperature $T=H/\pi$.Volovik2009 ; Volovik2023c ; Volovik2024c ; Maxfield2022 The same temperature determines the other activation processes, which are energetically forbidden in the Minkowski spacetime, but are allowed in the de Sitter background, see also Refs. Maldacena2015 ; Reece2023 (Sec. II.2). In Sec. II.3 it is shown that the local temperature also determines the Hawking temperature of radiation from the cosmological horizon without using the Euclidean action.
@@
All that suggests that it is natural to consider the temperature $T=H/\pi$ as the local temperature of the de Sitter vacuum. The local temperature leads to the local entropy of the de Sitter thermal bath (Sec. II.5), which integrated over the Hubble volume reproduces the Gibbons-Hawking entropy of the cosmological horizon (Sec.II.6). The effect of thermal fluctuations is discussed in Sec. II.7.
@@
Sec. III is devoted to the multi-metric gravity, which can be considered as ensemble of the sub-Universes, each described by its own metric $g_{\mu\nu(n)}$ (or tetrads $e^{a}_{\mu(n)}$), by its own gravitational coupling $K_{n}$ and cosmological constant $\Lambda_{n}$. The heat exchange between the sub-Universes leads to their thermalization – the formation of the Universe in which the sub-Universes have the common Hubble parameter and thus the common temperature.
@@
In Sec. IV the local entropy of the de Sitter state allows to consider the thermodynamics of the Schwarzschild black hole, assuming that the black hole can be formed by relaxation of the gravastar (Sec. IV.3) with the de Sitter core. Here the interior of the gravastar is represented by the contracting de Sitter state (Sec. IV.2), which has negative Hubble parameter, and thus the negative temperature $T=H/\pi<0$ and negative entropy. Since the cosmological and black hole horizons cancels each other, the gravastar has zero entropy. In the process of relaxation of the gravastar to the black hole the de Sitter core with its negative entropy shrinks, which results in the Hawking-Bekenstein entropy of the black hole horizon (Sec. IV.4). The entropy of the white hole is considered in Sec. IV.5, and the heat exchange between black holes in the multi-metric gravity is in Sec. IV.6.
@@
Sec. V is devoted to thermodynamics of de Sitter in the $f({\cal R})$ gravity. It contains the pair of the thermodynamically conjugate variables, $K$ and ${\cal R}$ (Sec. V.1). These variables together with the local temperature and local entropy provides the Gibbs-Duhem relation (Sec. V.2). The general result for the total entropy of the Hubble volume, $S_{\rm bulk}=sV_{H}=4\pi KA=S_{\rm hor}$ is obtained in Sec. V.3. This is considered on example of quadratic gravity in Sec. V.4.
@@
The Sec. VI is devoted to the de Sitter decay due to the thermalization of matter by the de Sitter heat bath, and finally the conclusion is in Sec. VII.
@@ II.1 Atom in de Sitter environment as thermometer
Now let us consider an atom at the origin, $r=0$. The atom is the external object in the de Sitter spacetime, which is playing the role of the detector (or the role of the static observer) in this spacetime. The electron bounded to an atom may absorb the energy from the gravitational field of the de Sitter background and escape from the electric potential barrier. If the ionization potential is much smaller than the electron mass but is much larger than the Hubble parameter, $H\ll\epsilon_{0}\ll m$, one can use the nonrelativistic quantum mechanics to estimate the tunneling rate through the barrier. The corresponding radial trajectory $p_{r}(r)$ is obtained from the classical equation $p_{r}^{2}/2m+p_{r}v(r)=-\epsilon_{0}$, where $p_{r}(r)v(r)$ is the Doppler shift:
@@
This is equivalent to the thermal thermal radiation with temperature $T=H/\pi$.
@@ II.2 Decay of composite particles in de Sitter spacetime
$w\sim\exp{\left(-\frac{\pi(m_{1}+m_{2}-m)}{H}\right)}\,.$ (5)
@@ II.3 Connection between the local and Hawking temperatures
The local temperature $T=H/\pi$ also determines the process of the Hawking radiation from the cosmological horizon and the Gibbons-Hawking temperature $T_{\rm GH}=T/2=H/2\pi$. The reason is that in the Hawking process, two particles are coherently created (the analog of cotunneling): one particle is created inside the horizon, while its partner is simultaneously created outside the horizon. The rate of the coherent radiation of two particles, each with energy $E$, is $w\propto\exp(-\frac{2E}{T})$. However, the observer can detect only the particle created inside the horizon. For this observer the creation rate $w\propto\exp(-\frac{2E}{T})$ is perceived as $w\propto\exp(-\frac{E}{T/2})=\exp(-\frac{E}{T_{\rm GH}})$ with the Gibbons-Hawking temperature $T_{\rm GH}=T/2=H/2\pi$.
@@
II.4 Accelerating detector: is there connection between the local process and Unruh radiation?
One may consider two independent processes related to the accelerating detector. One of them is the Unruh radiation, which could be linked to this apparent event horizon in the Rindler spacetime and to the corresponding Unruh temperature $T_{U}=a/2\pi$, where $a$ is acceleration.Unruh1976 ; Unruh2019 ; Unruh2022 The second one is the local process of ionization of the accelerating atom, which has no relation to the event horizon.Volovik2022e Similar to Eq.(4), the trajectory $p_{x}(x)$ of the radiated electron is obtained from the classical equation
@@ II.4 Accelerating detector: is there connection between the local process and Unruh radiation?
In principle, one can introduce the effective temperature, which is similar to the effective temperature $T_{\rm eff}\sim El$ of hopping electrons in a strong electric field $E$, where $l$ is the localization length.Shklovskii1992 ; Shklovskii2024 In our case, the analog of the localization length is $l=1/\sqrt{m\epsilon_{0}}$. But otherwise the local process of ionization of the accelerated atom is certainly non-thermal. Since $\epsilon_{0}\ll m$, the rate of ionization in Eq.(7) essentially exceeds the ionization rate in any thermal process, which involves acceleration, $w_{\rm thermal}\sim\exp(-\gamma\epsilon_{0}/a)$. This is very different from the ionization of the atom in the de Sitter environment, which looks thermal. This demonstrates the uniqueness of the de Sitter spacetime in producing the thermal bath with local temperature.
II.5 From local temperature to local entropy
Since the de Sitter state serves as the thermal bath for matter, it is not excluded that the de Sitter quantum vacuum may have its own temperature and entropy.Padmanabhan2020 If so, then the quasi-equilibrium states of the expanding Universe is described by two different temperatures: the temperature of the gravitational vacuum and the temperature of the matter degrees of freedom.Vergeles2023 In this section we discuss the pure de Sitter vacuum without the excited matter ignoring for the moment the thermal activated creation of matter from the vacuum. The excitation and thermalization of matter by the de Sitter thermal bath will be discussed in Sec. VI.
@@ II.5 From local temperature to local entropy
II.6 Hubble volume entropy vs entropy of the cosmological horizon
@@
where $A$ is the horizon area. This Hubble-volume entropy coincides with the Gibbons-Hawking entropy of the cosmological horizon. However, here it is the thermodynamic entropy coming from the local entropy of the de Sitter quantum vacuum, rather than the entropy of the horizon degrees of freedom.
Anyway, the relation between the bulk and surface entropies in the local vacuum thermodynamics suggests some holographic origin. Such bulk-surface correspondence is valid only in the ($3+1$)-dimension. In the general $d+1$ dimension of spacetime, the same approach gives the factor $(d-1)/2$ in the relation between the entropy of the Hubble volume and the Gibbons-Hawking entropy of the cosmological horizon, $S_{\rm bulk}=\frac{d-1}{2}\,S_{\rm hor}$.
This may add to the peculiarities of the $d=3$ space dimension,Barrow1983 where in particular the mass dimension of the gravitational coupling, $[K]=d-1$, coincides with the mass dimension of curvature, $[{\cal R}]=2$. The same concerns such pair of thermodynamically conjugate variables as electric field with the mass dimension $[E]=2$ and the electric induction with mass dimension $[D]=d-1$. Their dimensions also coincide only for $d=3$. Discussion of the natural dimensions of physical quantities in $d=3$ can be found in Ref. Volovik2022f .
II.7 Hubble volume vs the volume of Universe, and thermal fluctuations of de Sitter state

### v15

@@ I Introduction
The $f({\cal R})$ theory demonstrates that the effective gravitational coupling $K$ (it is the inverse Newton constant, $K=1/16\pi G$) and the scalar curvature ${\cal R}$ are connected by equation $K=df/d{\cal R}$. This suggests that $K$ and ${\cal R}$ are the thermodynamically conjugate variables.Volovik2023a ; Volovik2023b This pair of the gravitational variables is similar to the pair of the electrodynamic variables, electric field ${\bf E}$ and electric induction ${\bf D}$, which participate in the thermodynamics of dielectrics, and to the pair of magnetic thermodynamic variables, ${\bf B}$ and ${\bf H}$. Using the local thermodynamics with $T=H/\pi$, we obtained the general result for the total entropy of the Hubble volume, $S_{\rm bulk}=sV_{H}=4\pi KA=S_{\rm hor}$, where $K=df/d{\cal R}$ is the effective gravitational coupling and ${\cal R}=-12H^{2}$.
@@
In Sec. II we consider the elements of the local thermodynamics of the de Sitter state. It is based on the observation that the de Sitter vacuum serves as the thermal bath for the external matter – the matter immersed into the de Sitter environment. This is different from the traditional consideration of the vacua of the quantum fields in the de Sitter spacetime, which uses the Euclidean action method. Example of the influence of the de Sitter vacuum to the external matter is provided by an atom in the de Sitter environment (Sec. II.1). As distinct from the atom in the flat space, the atom in the de Sitter vacuum has a certain probability of ionization. The rate of ionization is similar to the rate of ionization in the presence of the thermal bath with temperature $T=H/\pi$.Volovik2009a ; Volovik2009 ; Volovik2023c ; Volovik2024c ; Maxfield2022 The same temperature determines the other activation processes, which are energetically forbidden in the Minkowski spacetime, but are allowed in the de Sitter background, see also Refs. Maldacena2015 ; Reece2023 (Sec. II.2). In Sec. II.3 it is shown that the local temperature also determines the Hawking temperature of radiation from the cosmological horizon without using the Euclidean action.
@@
All that suggests that it is natural to consider the temperature $T=H/\pi$ as the local temperature of the de Sitter vacuum. The local temperature leads to the local entropy of the de Sitter thermal bath (Sec. II.8), which integrated over the Hubble volume reproduces the Gibbons-Hawking entropy of the cosmological horizon (Sec.II.9). The effect of thermal fluctuations is discussed in Sec. III.1.
@@
Sec. IV is devoted to the multi-metric gravity, which can be considered as ensemble of the sub-Universes, each described by its own metric $g_{\mu\nu(n)}$ (or tetrads $e^{a}_{\mu(n)}$), by its own gravitational coupling $K_{n}$ and cosmological constant $\Lambda_{n}$. The heat exchange between the sub-Universes leads to their thermalization – the formation of the Universe in which the sub-Universes have the common Hubble parameter and thus the common temperature.
@@
In Sec. V the local entropy of the de Sitter state allows to consider the thermodynamics of the Schwarzschild black hole, assuming that the black hole can be formed by relaxation of the gravastar (Sec. V.3) with the de Sitter core. Here the interior of the gravastar is represented by the contracting de Sitter state (Sec. V.2), which has negative Hubble parameter, and thus the negative temperature $T=H/\pi<0$ and negative entropy. Since the cosmological and black hole horizons cancels each other, the gravastar has zero entropy. In the process of relaxation of the gravastar to the black hole the de Sitter core with its negative entropy shrinks, which results in the Hawking-Bekenstein entropy of the black hole horizon (Sec. V.4). The entropy of the white hole is considered in Sec. V.5, and the heat exchange between black holes in the multi-metric gravity is in Sec. V.6.
@@
Sec. VI is devoted to thermodynamics of de Sitter in the $f({\cal R})$ gravity. It contains the pair of the thermodynamically conjugate variables, $K$ and ${\cal R}$ (Sec. VI.1). These variables together with the local temperature and local entropy provides the Gibbs-Duhem relation (Sec. VI.2). The general result for the total entropy of the Hubble volume, $S_{\rm bulk}=sV_{H}=4\pi KA=S_{\rm hor}$ is obtained in Sec. VI.3. This is considered on example of quadratic gravity in Sec. VI.4.
@@
The Sec. VII is devoted to the de Sitter decay due to the thermalization of matter by the de Sitter heat bath, and finally the conclusion is in Sec. VIII.
@@ II.1 Atom in de Sitter environment as thermometer
Now let us consider an atom at the origin, $r=0$. The atom is the external object in the de Sitter spacetime, which is playing the role of the detector (or the role of the static observer) in this spacetime. The electron bounded to an atom may absorb the energy from the gravitational field of the de Sitter background and escape from the electric potential barrier. If the ionization potential is much smaller than the electron mass but is much larger than the Hubble parameter, $H\ll\epsilon_{0}\ll m$, one can use the nonrelativistic quantum mechanics to estimate the tunneling rate through the barrier. The corresponding radial trajectory $p_{r}(r)$ is obtained from the classical equation
$\frac{p_{r}^{2}}{2m}+p_{r}v(r)=-\epsilon_{0}\,,$ (2)
where $p_{r}(r)v(r)$ is the Doppler shift:
@@
This is equivalent to the thermal thermal radiation with temperature $T=H/\pi$, which is twice the Gibbons-Hawking temperature of the cosmological horizon, $T=2T_{\rm GH}$.
@@ II.2 Decay of composite particles in de Sitter spacetime
$w\sim\exp{\left(-\frac{\pi(m_{1}+m_{2}-m)}{H}\right)}=\exp{\left(-\frac{m_{1}+m_{2}-m}{T}\right)}\,.$ (6)
In the same way, the particle with mass $m$ can radiate the pair of the other particles moving in the opposite directions with the rate
$w\sim\exp{\left(-\frac{\pi(3m-m)}{H}\right)}=\exp{\left(-\frac{2m}{T}\right)}\,.$ (7)
For each of the two detectors, which detect the created particles moving in the opposite directions, the radiation rate looks as thermal with the Gibbons-Hawking temperature $T_{\rm GH}=T/2$. However, since two particles are created coherently, the temperature describing the radiation of the pair is twice the Gibbons-Hawking temperature. The same connection between the local temperature and the temperature of the Hawking radiation from the cosmological horizon is discussed in Sec. II.3.
@@
Note that the creation of new particles by the particle in the de Sitter environment demonstrates that the de Sitter vacuum is unstable towards the creation of matter. The consequence of such instability will be discussed in Sec. VII.
@@ II.3 Connection between the local and Hawking temperatures
The local temperature $T=H/\pi$ also determines the process of the Hawking radiation from the cosmological horizon and the Gibbons-Hawking temperature $T_{\rm GH}=T/2=H/2\pi$. The reason is that in the Hawking process, two particles are coherently created (the analog of cotunneling): one particle is created inside the horizon, while its partner is simultaneously created outside the horizon.Parikh2002 The rate of the coherent radiation of two particles, each with energy $E$, is $w\propto\exp(-\frac{2E}{T})$. However, the observer who uses the Unruh-DeWitt detector can detect only the particle created inside the horizon. For this observer the creation rate $w\propto\exp(-\frac{2E}{T})$ is perceived as
$w\propto\exp\left(-\frac{E}{T/2}\right)=\exp\left(-\frac{E}{T_{\rm GH}}\right)\,,$ (8)
with the Gibbons-Hawking temperature $T_{\rm GH}=T/2=H/2\pi$.
@@
II.4 Two detectors: exited atom vs ionized atom
Note the main difference between the temperature measured by the observer using the Unruh-DeWitt detector (see e.g. Ref.Conroy2022 and references therein) and the temperature measured by the observer using ionization of an atom. The ionization process is possible, because the radiated electron moves far away from the atom to position $r=r_{0}$, where its negative gravitational energy compensates the ionization potential. In the Unruh-DeWitt detector, which corresponds to the two-level atom interacting with a quantum field, the electron in the atom is excited but remains in the same position in the same atom. That is why such excitation of electron may only come either by the Hawking radiation, or by radiation of photons, see Sec. II.5.
However, in the de Sitter case there is no difference in temperatures measured by two detectors. If in the Unruh-DeWitt detector experiments the Hawking temperature is properly interpreted according to Eq.(8), the measured temperature of the Hawking radiation is also $T=2T_{\rm GH}$. So, in spite of different physical principles, both detectors in the de Sitter environment show the same physical temperature, $T=2T_{\rm GH}$. This also demonstrates the uniqueness of the de Sitter state.
II.5 Radiation of photons by atom in de Sitter environment
Instead of the radiation of electron from the atom, one can consider the radiation of photon by the same atom. In both cases the escape of photon or electron from the atom provides the negative gravitational energy, which allows to consider the processes which are prohibited in the Minkowski state. Let us consider the atom in the excited state with energy $E_{0}+\epsilon$, where $E_{0}$ is the ground state energy. In the de Sitter environment, the radiation of photon is possible even if its energy $cp$ exceeds the energy difference between the excited and the ground state level, i.e. when $cp>\epsilon$.
For the relativistic photon the equation (2) takes the form (here we use $c=1$):
$\sqrt{p_{r}^{2}+p_{\perp}^{2}}+p_{r}v(r)=\epsilon\,.$ (9)
This gives the trajectory of photon
$p_{r}(r)=\frac{1}{1-H^{2}r^{2}}\left[-\epsilon Hr\pm\sqrt{\epsilon^{2}+p_{\perp}^{2}(H^{2}r^{2}-1)}\right]\,.$ (10)
For $cp_{\perp}>\epsilon$, the integration over the classically forbidden region, $r<r_{0}$, where $H^{2}r_{0}^{2}=1-\epsilon^{2}/p_{\perp}^{2}$, gives the following radiation rate of photon by excited atom:
$w\sim\exp{\left(-2\,{\rm Im}\,S\right)}=\exp\left(-\frac{\pi(cp_{\perp}-\epsilon)}{H}\right)\,\,,\,\,cp_{\perp}>\epsilon\,.$ (11)
Since $r_{0}<1/H$, the trajectory of photon in this process is inside the cosmological horizon, and the radiation rate is again determined by the local temperature $T=H/\pi$.
Eq.(11) demonstrates, that radiation takes place even if the atom is in the ground state, i.e. when $\epsilon\rightarrow 0$. In this case Eq.(11) describes the process of radiation of photon in the de Sitter environment:
$w\sim\exp\left(-\frac{cp}{T}\right)\,\,,\,\,T=\frac{H}{\pi}\,.$ (12)
In this process, the atom in its ground state serves as an external object, which violates the de Sitter symmetry and thus provides the nonzero matrix element for the radiation of photon.
II.6 Accelerating detector: is there connection between the local process and Unruh radiation?
In the same way as in Sections II.4 and II.5, one may consider different independent processes related to the accelerating detector. One of them is the Unruh radiation measured by the Unruh-DeWitt detector, which could be linked to the apparent event horizon in the Rindler spacetime and to the corresponding Unruh temperature $T_{U}=a/2\pi$, where $a$ is acceleration.Unruh1976 ; Unruh2019 ; Unruh2022 The other processes, such as the local process of ionization of the accelerating atom,Volovik2022e have no relation to the event horizon.
Let us consider the process of ionization of atom. Similar to Eq.(5), the trajectory $p_{x}(x)$ of the radiated electron is obtained from the classical equation
@@ II.6 Accelerating detector: is there connection between the local process and Unruh radiation?
In principle, one can introduce the effective temperature, which is similar to the effective temperature $T_{\rm eff}\sim El$ of hopping electrons in a strong electric field $E$, where $l$ is the localization length.Shklovskii1992 ; Shklovskii2024 In our case, the analog of the localization length is $l=1/\sqrt{m\epsilon_{0}}$. But otherwise the local process of ionization of the accelerated atom is certainly non-thermal. Since $\epsilon_{0}\ll m$, the rate of ionization in Eq.(14) essentially exceeds the ionization rate in any thermal process, which involves acceleration, $w_{\rm thermal}\sim\exp(-\gamma\epsilon_{0}/a)$. This is very different from the ionization of the atom in the de Sitter environment, which looks thermal.
In principle, with the same approach one can also calculate the Unruh radiation of photons using different arrangements of detectors. The state of the art in theoretical understanding of Unruh radiation can be found in Refs. Gregori2023 ; Gregori2022 ; Matsuda2024 and references therein. Our examples demonstrate that the rate of Unruh radiation depends on details of the considered processes and it is not necessarily thermal.
II.7 de Sitter symmetry and de Sitter heat bath
As distinct from the Unruh effect, different arrangements of detectors in the de Sitter environment show the same temperature. This demonstrates the uniqueness of the de Sitter spacetime in producing the thermal bath with local temperature. The reason for that is that the de Sitter spacetime is homogeneous under the combination of translation and the proper conformal transformations.Aldrovandi2007 ; Kamenshchik2008 In the PG metric it is the invariance of the de Sitter state under the modified translations, ${\bf r}\rightarrow{\bf r}-e^{Ht}{\bf a}$, which for $H\rightarrow 0$ corresponds to the conventional invariance under translations ${\bf r}\rightarrow{\bf r}-{\bf a}$ in Minkowski spacetime. Due to this combined translational symmetry, all the comoving observers at any point of the de Sitter space observe the same temperature, $T=H/\pi$. That is why one may conclude, that the de Sitter state is the heat bath produced by gravity.
The uniqueness of the de Sitter thermal state lies in the fact that its temperature does not violate the de Sitter symmetry, and thus does not require the preferred reference frame. This is distinct from thermal state of matter, which always has a preferred reference frame where matter is at rest. As a result, thermal matter violates the de Sitter symmetry, which leads to the heat exchange between the two thermal subsystems, gravity and matter.
II.8 From local temperature to local entropy
The gravity subsystem – de Sitter quantum vacuum – has its own temperature and entropy, see also Ref.Padmanabhan2020 . On the other hand the de Sitter spacetime serves as the thermal bath for matter. Then the quasi-equilibrium states of the expanding Universe can be described by two different temperatures: the temperature of the gravitational vacuum (temperature of dark energy) and the temperature of the matter degrees of freedom.Vergeles2023 In this section we discuss the pure de Sitter vacuum without the excited matter ignoring for the moment the thermal activated creation of matter from the vacuum. The excitation and thermalization of matter by the de Sitter thermal bath will be discussed in Sec. VII.
@@ II.8 From local temperature to local entropy
The equation (16) demonstrates that the thermal properties of the de Sitter state are similar to that of the non-relativistic Fermi liquid, where the entropy density is also linear in temperature:
$s_{\rm FL}=\frac{p_{F}^{2}}{3v_{F}}T\,.$ (17)
The Fermi velocity $v_{F}$ and Fermi momentum $p_{F}$ of this cosmological analog of Fermi liquid are on the order of the speed of light and the inverse Planck length correspondingly, $v_{F}\sim c$ and $p_{F}\sim 1/l_{\rm Pl}$. Of course, this is only the parallel, and one should not identify the de Sitter vacuum with the real Fermi liquid. However, the real Fermi surface may form inside the black hole horizon.HuhtalaVolovik2002 ; Zubkov2019
II.9 Hubble volume entropy vs entropy of the cosmological horizon
@@
where $A$ is the horizon area.
The Hubble-volume entropy coincides with the Gibbons-Hawking entropy of the cosmological horizon. However, here it is the thermodynamic entropy coming from the local entropy of the de Sitter quantum vacuum, rather than the entropy of the horizon degrees of freedom. It is typically assumed that the horizon microstates are concentrated in the region of the Planck length $l_{\rm Pl}=1/M_{\rm Pl}$ in the vicinity of horizon, and this leads to the area law for the total entropy inside the horizon.Padmanabhan2010 The local thermodynamics leads to the same area law.
The connection between the bulk entropy of the region inside the horizon due to the local vacuum thermodynamics and the surface entropy of the horizon has the holographic origin, which is related to the horizon as null surface.Padmanabhan2010 This is the reason why the presence of the horizon allows to know the total entropy of the Hubble volume without knowledge on the local entropy inside the horizon.
This bulk-surface correspondence in Eq.(18) is valid only in the ($3+1$)-dimension. In the general $d+1$ dimension of spacetime, the same approach gives the factor $(d-1)/2$ in the relation between the entropy of the Hubble volume and the Gibbons-Hawking entropy of the cosmological horizon, $S_{\rm bulk}=\frac{d-1}{2}\,S_{\rm hor}$. This may add to the peculiarities of the $d=3$ space dimension,Barrow1983 where in particular the mass dimension of the gravitational coupling, $[K]=d-1$, coincides with the mass dimension of curvature, $[{\cal R}]=2$. The same concerns such pair of thermodynamically conjugate variables as electric field with the mass dimension $[E]=2$ and the electric induction with mass dimension $[D]=d-1$. Their dimensions also coincide only for $d=3$. Discussion of the natural dimensions of physical quantities in $d=3$ can be found in Ref. Volovik2022f .
III Hubble volume vs the volume of Universe
III.1 Thermal fluctuations of de Sitter state